trinitycook489:

What is the first step in solving the equation ? 2/3 x+ 1/3 x+2=5

umm:

Does your equation look like this?\[\frac{ 2 }{ 3 }x+\frac{ 1 }{ 3 }x+2=5\]

trinitycook489:

yes!

umm:

Perfect. Is there answer choices for this?

trinitycook489:

yes there are A.) Subtract One-third from each side of the equation. B.) Add 2 to each side of the equation. C.) Combine like terms. D.) Multiply each side of the equation by 5.

umm:

Ah, we'll let's go over it step-by-step, and we can use process of elimination.

trinitycook489:

ok but i only got 8 mins so can we hurry

umm:

Is this a test?

trinitycook489:

nobut i got to go in 8 minutes

umm:

Oh, OK. One second.

trinitycook489:

ok

trinitycook489:

3 mins left

trinitycook489:

2 mins

trinitycook489:

1 minute

umm:

First thing's first, we'd simplify both sides of the equation.\[\frac{ 2 }{ 3 }x+\frac{ 1 }{ 3 }x+2=5\] And then we combine like terms \[(\frac{ 2 }{ 3 }x+\frac{ 1 }{ 3 })+ (2)=5\]

umm:

You should be able to see the first step here.

trinitycook489:

k thx byee now

umm:

Good luck.